OpenStudy (anonymous):

find the intervals of increase and decrease for e^2x-e^x, please help!

OpenStudy (michele_laino):

here we have to compute the first derivative of your function

OpenStudy (michele_laino):

I got this: \[\Large f'\left( x \right) = {e^x}\left( {2{e^x} - 1} \right)\]

OpenStudy (michele_laino):

now the intervals of the real line in which f(x) is increasing, are given by the subsequent condition: \[\Large f'\left( x \right) > 0 \Rightarrow {e^x}\left( {2{e^x} - 1} \right) > 0\]

OpenStudy (michele_laino):

which, since e^x is always positive, is equivalent to this one: \[\Large 2{e^x} - 1 > 0\] please solve that inequality, what do you get?

OpenStudy (anonymous):

e^x>1/2

OpenStudy (michele_laino):

please now solve it for x

OpenStudy (anonymous):

would it be ln(e)>1/2

OpenStudy (michele_laino):

no, since if we take the logarithm of both sides, we get: \[\Large \begin{gathered} x\left( {\ln e} \right) > \ln \left( {1/2} \right) \hfill \\ x > - \ln 2 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

am I right?

OpenStudy (michele_laino):

please remember that: \[\Large \begin{gathered} \ln e = 1 \hfill \\ \ln \left( {1/2} \right) = \ln 1 - \ln 2 = 0 - \ln 2 = - \ln 2 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

oh okay i see it now

OpenStudy (michele_laino):

so your function is increasing in: [-ln 2, +infinity) and decreasing, in: (-infinity, -ln2)

OpenStudy (michele_laino):

the domain of your function is all the real line

OpenStudy (anonymous):

okay! and to find the local max and min i would have to find the second derivative and plug in -ln2 for x right?

OpenStudy (michele_laino):

no, the local max and min comes from the first derivative

OpenStudy (michele_laino):

now we have this: